Can you please help me determine what would be the next step in finding a solution to this equation? 8x = 6 + 2x
Answer by rchill(405)

Subtract 2x from both sides: 8x-2x=6+2x-2x, which simplifies to 6x=6. Now divide both sides by 6 and get x=1.